OpenStudy (anonymous):

Two jets leave Denver at 9:00 a.m. , one flying east at a Soren 50 km/h greater than the other, which is traveling west. At 11:00 a.m. The planes are 2500 km apart. Find their speeds

OpenStudy (anonymous):

West-bound DATA: time = 2 hrs ; rate = x km/h ; distance = rt = 2x km ----------------------- East bound DATA: time = 2 hrs ; rate = (x+50)km/h ; distance = 2(x+50) km ----------------------- Equation: distance + distance = 2500 2x + 2(x+50) = 2500 x + x+50 = 1250 2x = 1200 x = 600 km/h (speed of the West-bound plane) --- x+50 = 650 km/h (speed of the East-bound plane)